OpenStudy (anonymous):

Efficacy of a drug. Seventy-five (75) people with hypertension were given a new drug for the treatment of their condition. Blood pressure was lowered sufficiently to reduce the risk of serious outcomes in 20 of the 75 patients. (A) In what proportion of the sample did the drug work? (B) Calculate a 95% confidence interval for the proportion. Show all work, and interpret your findings. (C) How large a sample would be needed to reduce the margin of error to .05?

OpenStudy (anonymous):

Binomial distribution problem. p = (20/75), q = 1-p, n=75, r=20

OpenStudy (anonymous):

How do I proceed in solving it?

OpenStudy (anonymous):

A. is p B. review binomial distribution, which has mean = np and standard deviation of sqrt(npq). Divide these by n to get the proportion and its standard deviation. Check text to find confidence interval limits. Assume "margin of error" means the width of the confidence interval, and see ow large n must be to get this to 5%. Must go.